# Parabolic Equations with Singular Coefficients and Boundary Data: Analysis and Numerical Simulations

### v1

@@ 4. Very weak solutions
Since equation (4.2) involves products of distributions, we replace it with a regularised problem using smooth approximations of all data and coefficients.
Let $(q_{\varepsilon})$ be an $L^{\infty}$-moderate regularisation of $q$, $(a_{\varepsilon})$ a $W^{1,\infty}$-moderate regularisation of $a$, with $a_{\varepsilon}\geq\alpha>0$, $(b_{\varepsilon})$ a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation of $b$, $(u_{0,\varepsilon})$ an $H^{1}$-moderate regularisation of $u_{0}$, $(g_{0,\varepsilon}),(g_{1,\varepsilon})$ time-distributional moderate regularisations of $g_{0},g_{1}$, and $(f_{\varepsilon})$ an $L^{2}$-moderate regularisation of $f$. Then, for each $\varepsilon\in(0,1]$, consider the regularised problem (4.3).
For each fixed $\varepsilon\in(0,1]$, the coefficients and data in (4.3) are smooth, and the diffusion coefficient $a_{\varepsilon}$ is uniformly elliptic. Hence, by classical parabolic theory and the energy estimate (see Corollary 3.2), there exists a unique weak solution
$u_{\varepsilon}\in H^{1}([0,T];L^{2}(0,1))\cap L^{2}([0,T];H^{2}(0,1)).$
Moreover, Corollary 3.2 implies the energy estimate
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\qquad\times\left(\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\right).$ (4.4)
Using the moderateness assumptions for all regularised nets, the right-hand side is bounded by $C\varepsilon^{-N}$ for some $N\in\mathbb{N}$, uniformly in $\varepsilon\in(0,1]$. Thus,
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N},$
so $(u_{\varepsilon})$ is moderate in $H^{1}L^{2}\cap L^{2}H^{2}$, and therefore, by Definition 6, $(u_{\varepsilon})$ is a very weak solution to (4.2). ∎
@@
Let $(u_{\varepsilon})$, $(\tilde{u}_{\varepsilon})$ be two nets in $L^{\infty}([0,T];L^{2}(0,1))$. We say that the net $(u_{\varepsilon}-\tilde{u}_{\varepsilon})$ is $L^{2}$-negligible if, for every $N\in\mathbb{N}$ there exist a constant $C>0$ such that
$\|u_{\varepsilon}(t,\cdot)-\tilde{u}_{\varepsilon}(t,\cdot)\|_{L^{2}}\leq C\varepsilon^{N},$
@@
$\left\{\begin{array}[]{l}\partial_{t}\tilde{u}_{\varepsilon}(t,x)-\partial_{x}(\tilde{a}_{\varepsilon}(t,x)\partial_{x}\tilde{u}_{\varepsilon})(t,x)+\tilde{b}_{\varepsilon}(t,x)\partial_{x}\tilde{u}_{\varepsilon}(t,x)+\tilde{q}_{\varepsilon}(x)\tilde{u}_{\varepsilon}(t,x)\\ \hfill=\tilde{f}_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ \tilde{u}_{\varepsilon}(0,x)=\tilde{u}_{0,\varepsilon}(x),\,\,\,x\in(0,1),\\ \tilde{u}_{\varepsilon}(t,0)=\tilde{g}_{0,\varepsilon}(t),\quad\tilde{u}_{\varepsilon}(t,1)=\tilde{g}_{1,\varepsilon}(t),\quad t\in[0,T].\end{array}\right$. (4.6)
Definition 8 (Uniqueness of the very weak solution).
Let $a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $q\in\mathcal{D}^{\prime}(0,1)$, $g_{0},\,g_{1}\in\mathcal{D}^{\prime}([0,T])$. We say that the initial/boundary problem (4.2) has a unique very weak solution, if for all $W^{1,\infty}(L^{2})$, log-$W^{1,\infty}$-moderate nets $a_{\varepsilon}$, $\tilde{a}_{\varepsilon}$, $b_{\varepsilon}$, $\tilde{b_{\varepsilon}}$, such that $(a_{\varepsilon}-\tilde{a}_{\varepsilon})$, $(b_{\varepsilon}-\tilde{b}_{\varepsilon})$ are $W^{1,\infty}$-negligible and log-$L^{\infty}(W^{1,\infty})$-negligible; $L^{\infty}$-moderate nets $q_{\varepsilon}$, $\tilde{q}_{\varepsilon}$, such that $(q_{\varepsilon}-\tilde{q}_{\varepsilon})$ are $L^{\infty}$-negligible; time-distributional moderate nets $g_{0,\varepsilon},\,g_{1,\varepsilon},\,\tilde{g}_{0,\varepsilon},\,\tilde{g}_{1,\varepsilon}$, such that $(g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon})$, $(g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon})$ are $H^{1}$-negligible; for all $L^{2}$-moderate nets $f_{\varepsilon},\,\tilde{f}_{\varepsilon}$, such that $(f_{\varepsilon}-\tilde{f}_{\varepsilon})$ are $L^{2}$-negligible; for all $H^{1}$-moderate nets $u_{0,\varepsilon},\,\tilde{u}_{0,\varepsilon}$ such that $(u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon})$ is $H^{1}$-negligible, we have that $u_{\varepsilon}-\tilde{u}_{\varepsilon}$ is $H^{1}(L^{2})\cap L^{2}(H^{2})$-negligible.
@@
Let us denote by $u_{\varepsilon}$ and $\tilde{u}_{\varepsilon}$ the families of solutions for the initial/boundary problems (4.5) and (4.6) correspondingly. Let $U_{\varepsilon}$ represents the difference between these nets $U_{\varepsilon}(t,\cdot):=u_{\varepsilon}(t,\cdot)-\tilde{u}_{\varepsilon}(t,\cdot)$ then $U_{\varepsilon}$ solves the equation
$\left\{\begin{array}[]{l}\partial_{t}U_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}U_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}U_{\varepsilon}(t,x)+q_{\varepsilon}(x)U_{\varepsilon}(t,x)\\ \hfill=F_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ U_{\varepsilon}(0,x)=(u_{0,\varepsilon}-\tilde{u}_{0,\varepsilon})(x),\,\,\,x\in(0,1),\\ U_{\varepsilon}(t,0)=(g_{0,\varepsilon}-\tilde{g}_{0,\varepsilon})(t),\,U_{\varepsilon}(t,1)=(g_{1,\varepsilon}-\tilde{g}_{1,\varepsilon})(t),\,\,t\in[0,T],\end{array}\right$. (4.7)
@@
$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\left(\big(\tilde{a}_{\varepsilon}(t,x)-a_{\varepsilon}(t,x)\big)\partial_{x}\tilde{u}_{\varepsilon}\right)(t,x)+\big(\tilde{b}_{\varepsilon}(t,x)-b_{\varepsilon}(t,x)\big)\partial_{x}\tilde{u}_{\varepsilon}(t,x)$
$+\big(\tilde{q}_{\varepsilon}(x)-q_{\varepsilon}(x)\big)\tilde{u}_{\varepsilon}(t,x)+f_{\varepsilon}(t,x)-\tilde{f}_{\varepsilon}(t,x)$
@@
Taking the $H^{1}(L^{2})\cap L^{2}(H^{2})$-norm of the $U_{\varepsilon}$ by using (4), we obtain
@@
$\qquad\leq\frac{C}{\alpha}\exp\left\{T\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right\}\Big(\|a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{2}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\quad+\frac{1}{\sqrt{\alpha}}\left(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\right)\Big)$
[…]
$\quad\times\left(\|U_{\varepsilon}(0,\cdot)\|_{L^{2}}+\|F_{\varepsilon}\|_{L^{2}}+\|U_{\varepsilon}(\cdot,0)\|_{H^{1}}+\|U_{\varepsilon}(\cdot,1)\|_{H^{1}}\right)\Bigg).$
@@
$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\partial_{x}((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon})\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|(\tilde{b}_{\varepsilon}-b_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|(\tilde{q}_{\varepsilon}-q_{\varepsilon})\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$ (4.8)
@@
$\partial_{x}\big((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}\big)=\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon}+(\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}^{2}\tilde{u}_{\varepsilon}.$
@@
$\|\partial_{x}\cdot((\tilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\tilde{u}_{\varepsilon})\|$ $_{L^{2}([0,T];L^{2}(0,1))}\leq\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\quad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
By Proposition 2.4 and Corollary 3.2, this quantity is finite and moderate. Then, we have
$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\tilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\quad+\|\partial_{x}(\tilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|\tilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|\tilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\tilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\tilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$

### v2

@@ 3. Very weak solutions for the non-homogeneous Dirichlet conditions
Since (3.1) involves products of distributions, we consider the regularised problem (3.2) for each $\varepsilon\in(0,1]$, where $a_{\varepsilon},b_{\varepsilon},q_{\varepsilon},f_{\varepsilon},u_{0,\varepsilon},g_{0,\varepsilon},g_{1,\varepsilon}$ are smooth regularisations satisfying the moderateness assumptions of Definition 6 and $a_{\varepsilon}\geq\alpha>0$.
For fixed $\varepsilon$, the coefficients are bounded and $a_{\varepsilon}$ is uniformly elliptic, hence there exists a unique weak solution $u_{\varepsilon}$ to (3.2). Moreover, Proposition 3.2 yields
$\partial_{t}u_{\varepsilon},\ \partial_{x}^{2}u_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$
together with the estimate (3.2).
We now show moderateness of $(u_{\varepsilon})$. By the assumptions on the regularisations, we have polynomial bounds for $\|u_{0,\varepsilon}\|_{H^{1}_{0}}$, $\|f_{\varepsilon}\|_{L^{2}}$, $\|a_{\varepsilon}\|_{W^{1,\infty}}$, $\|\partial_{t}a_{\varepsilon}\|_{L^{\infty}}$, $\|b_{\varepsilon}\|_{L^{\infty}}$, and log-bounds for $\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}}$ and $\|q_{-,\varepsilon}\|_{L^{\infty}}$. Consequently,
$\exp\!\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}}+2\|q_{-,\varepsilon}\|_{L^{\infty}}\right)\right\}\leq C\,\varepsilon^{-M}$
for some $M\in\mathbb{N}$, and all remaining factors on the right-hand side of (3.2) are bounded by $C\varepsilon^{-M}$ as well. Hence there exist $C>0$ and $N\in\mathbb{N}$ such that
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|\partial_{x}^{2}u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}\leq C\,\varepsilon^{-N},\qquad\varepsilon\in(0,1].$
This is exactly the moderateness condition in Definition 6. Therefore $(u_{\varepsilon})$ is a very weak solution to (3.1). ∎
@@
Let $(u_{\varepsilon})$, $(\widetilde{u}_{\varepsilon})$ be two nets in $L^{\infty}([0,T];L^{2}(0,1))$. We say that the net $(u_{\varepsilon}-\widetilde{u}_{\varepsilon})$ is $L^{\infty}(L^{2})$-negligible if, for every $N\in\mathbb{N}$ there exist a constant $C>0$ such that
$\|u_{\varepsilon}(t,\cdot)-\widetilde{u}_{\varepsilon}(t,\cdot)\|_{L^{2}}\leq C\varepsilon^{N},$
@@
$\left\{\begin{array}[]{l}\partial_{t}\widetilde{u}_{\varepsilon}(t,x)-\partial_{x}(\widetilde{a}_{\varepsilon}(t,x)\partial_{x}\widetilde{u}_{\varepsilon})(t,x)+\widetilde{b}_{\varepsilon}(t,x)\partial_{x}\widetilde{u}_{\varepsilon}(t,x)+\widetilde{q}_{\varepsilon}(x)\widetilde{u}_{\varepsilon}(t,x)\\ \hfill=\widetilde{f}_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ \widetilde{u}_{\varepsilon}(0,x)=\widetilde{u}_{0,\varepsilon}(x),\,\,\,x\in(0,1),\\ \widetilde{u}_{\varepsilon}(t,0)=\widetilde{g}_{0,\varepsilon}(t),\quad\widetilde{u}_{\varepsilon}(t,1)=\widetilde{g}_{1,\varepsilon}(t),\quad t\in[0,T].\end{array}\right$. (3.11)
Definition 8 (Uniqueness of very weak solutions).
We say that the initial–boundary value problem (3.1) has a unique very weak solution if the following holds:
Let $(a_{\varepsilon},b_{\varepsilon},q_{\varepsilon},u_{0,\varepsilon},g_{0,\varepsilon},g_{1,\varepsilon},f_{\varepsilon})$ and $(\widetilde{a}_{\varepsilon},\widetilde{b}_{\varepsilon},\widetilde{q}_{\varepsilon},\widetilde{u}_{0,\varepsilon},\widetilde{g}_{0,\varepsilon},\widetilde{g}_{1,\varepsilon},\widetilde{f}_{\varepsilon})$ be two families of regularisations satisfying the assumptions of Definition 6, and let $u_{\varepsilon}$, $\widetilde{u}_{\varepsilon}$ be the corresponding solutions of (3.10) and (3.11).
Assume in addition that the differences of the regularisations are negligible in the corresponding norms, i.e.
$a_{\varepsilon}-\widetilde{a}_{\varepsilon}\ \text{is $W^{1,\infty}$-negligible},\qquad b_{\varepsilon}-\widetilde{b}_{\varepsilon}\ \text{is log-$L^{\infty}(W^{1,\infty})$-negligible},$
$q_{\varepsilon}-\widetilde{q}_{\varepsilon}\ \text{is $L^{\infty}$-negligible},\qquad u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon}\ \text{is $H^{1}$-negligible},$
$g_{j,\varepsilon}-\widetilde{g}_{j,\varepsilon}\ \text{is $H^{1}(0,T)$-negligible for }j=0,1,\qquad f_{\varepsilon}-\widetilde{f}_{\varepsilon}\ \text{is $L^{2}(0,T;L^{2})$-negligible}.$
Then
$u_{\varepsilon}-\widetilde{u}_{\varepsilon}\ \text{is $H^{1}(0,T;L^{2}(0,1))\cap L^{2}(0,T;H^{2}(0,1))$-negligible}.$
@@
Let us denote by $u_{\varepsilon}$ and $\widetilde{u}_{\varepsilon}$ the families of solutions for the initial/boundary problems (3.10) and (3.11) correspondingly. Let $U_{\varepsilon}$ represents the difference between these nets $U_{\varepsilon}(t,\cdot):=u_{\varepsilon}(t,\cdot)-\widetilde{u}_{\varepsilon}(t,\cdot)$ then $U_{\varepsilon}$ solves the equation
$\left\{\begin{array}[]{l}\partial_{t}U_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}\partial_{x}U_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}U_{\varepsilon}(t,x)+q_{\varepsilon}(x)U_{\varepsilon}(t,x)\\ \hfill=F_{\varepsilon}(t,x),\quad(t,x)\in[0,T]\times(0,1),\\ U_{\varepsilon}(0,x)=(u_{0,\varepsilon}-\widetilde{u}_{0,\varepsilon})(x),\,\,\,x\in(0,1),\\ U_{\varepsilon}(t,0)=(g_{0,\varepsilon}-\widetilde{g}_{0,\varepsilon})(t),\,U_{\varepsilon}(t,1)=(g_{1,\varepsilon}-\widetilde{g}_{1,\varepsilon})(t),\,\,t\in[0,T],\end{array}\right$. (3.12)
@@
$F_{\varepsilon}(t,x)$ $:=-\partial_{x}\left(\big(\widetilde{a}_{\varepsilon}(t,x)-a_{\varepsilon}(t,x)\big)\partial_{x}\widetilde{u}_{\varepsilon}\right)(t,x)+\big(\widetilde{b}_{\varepsilon}(t,x)-b_{\varepsilon}(t,x)\big)\partial_{x}\widetilde{u}_{\varepsilon}(t,x)$
$+\big(\widetilde{q}_{\varepsilon}(x)-q_{\varepsilon}(x)\big)\widetilde{u}_{\varepsilon}(t,x)+f_{\varepsilon}(t,x)-\widetilde{f}_{\varepsilon}(t,x)$
@@
Taking the $H^{1}(L^{2})\cap L^{2}(H^{2})$-norm of the $U_{\varepsilon}$ by using (3.2), we obtain
@@
$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
$\qquad\times\Big(\sqrt{\alpha}\|U_{\varepsilon}(0,\cdot)\|_{H^{1}_{0}(0,1)}+\|F_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|U_{\varepsilon}(\cdot,0)\|_{H^{1}([0,T])}+\|U_{\varepsilon}(\cdot,1)\|_{H^{1}([0,T])}\Big)$
$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|U_{\varepsilon}(\cdot,0)\|_{H^{1}([0,T])}+\|U_{\varepsilon}(\cdot,1)\|_{H^{1}([0,T])}\big)\Big).$
@@
$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\partial_{x}((\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon})\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|(\widetilde{b}_{\varepsilon}-b_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|(\widetilde{q}_{\varepsilon}-q_{\varepsilon})\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$ (3.13)
@@
$\partial_{x}\big((\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon}\big)=\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon}+(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}^{2}\widetilde{u}_{\varepsilon}.$
@@
$\|\partial_{x}\cdot((\widetilde{a}_{\varepsilon}-a_{\varepsilon})\partial_{x}\widetilde{u}_{\varepsilon})\|$ $_{L^{2}([0,T];L^{2}(0,1))}\leq\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\quad+\|\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
By Lemma 2.4 and Proposition 3.2, this quantity is finite and moderate. Then, we have
$\|F_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$ $\leq\|\widetilde{a}_{\varepsilon}-a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}^{2}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\quad+\|\partial_{x}(\widetilde{a}_{\varepsilon}-a_{\varepsilon})\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\,\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|\widetilde{b}_{\varepsilon}-b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\|\partial_{x}\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$+\|\widetilde{q}_{\varepsilon}-q_{\varepsilon}\|_{L^{\infty}(0,1)}\|\widetilde{u}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|f_{\varepsilon}-\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$